OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 18. What is the value of x when y = 72? A. 54 B. 12 C. 6 D. 8

OpenStudy (anonymous):

b.12

OpenStudy (amistre64):

maybe its a.54 how do we decide?

OpenStudy (anonymous):

Umm.. idk. lol Thanks you guys:)

OpenStudy (amistre64):

its the "idk" that is the problem :) just spouting out answers is no way to help a person study in my opinion

OpenStudy (anonymous):

x1/y1=x2/y2 3/18=x/72 x=72/6 x=12

OpenStudy (anonymous):

Thank you<3

OpenStudy (amistre64):

that seems better ;)